A bucket mass of 1 kg is whirled in a vertical circle of radius 1 m. How fast must the bucket move at the top of the circle so that the rope doesn't slack?

1
Expert's answer
2022-06-07T08:12:02-0400

"F_1=F_2,"

"mg=m\frac{v^2}r,"

"v=\sqrt{gr}=3.13~\frac ms."